OpenStudy (anonymous):

eilee is now 24 years older than her daughter. in 8 years, Eilee will be twice as old as her daughter will be then. how old is her daughter now?

OpenStudy (jagatuba):

E=D+24 (E+8)/2=D+8

OpenStudy (anonymous):

Oh thanks :)

OpenStudy (jagatuba):

Or E+8=2(D+8) if you prefer.

OpenStudy (jagatuba):

Can you solve from here or do you need more help?

OpenStudy (anonymous):

I can :D

OpenStudy (jagatuba):

Cool